# AutoAL: Automated Active Learning with Differentiable Query Strategy Search

### v1

@@ Abstract
As deep learning continues to evolve, the need for data efficiency becomes increasingly important. Considering labeling large datasets is both time-consuming and expensive, active learning (AL) provides a promising solution to this challenge by iteratively selecting the most informative subsets of examples to train deep neural networks, thereby reducing the labeling cost. However, the effectiveness of different AL algorithms can vary significantly across data scenarios, and determining which AL algorithm best fits a given task remains a challenging problem. This work presents the first differentiable AL strategy search method, named AutoAL, which is designed on top of existing AL sampling strategies. AutoAL consists of two neural nets, named SearchNet and FitNet, which are optimized concurrently under a differentiable bi-level optimization framework. For any given task, SearchNet and FitNet are iteratively co-optimized using the labeled data, learning how well a set of candidate AL algorithms perform on that task. With the optimal AL strategies identified, SearchNet selects a small subset from the unlabeled pool for querying their annotations, enabling efficient training of the task model. Experimental results demonstrate that AutoAL consistently achieves superior accuracy compared to all candidate AL algorithms and other selective AL approaches, showcasing its potential for adapting and integrating multiple existing AL methods across diverse tasks and domains. Code will be available at: https://github.com/haizailache999/AutoAL.
@@ 1 Introduction
With the development of deep learning (DL) techniques, large volumes of high-quality labeled data have become increasingly crucial as the model training processes are usually data hungry  Ren et al. (2021); Zhan et al. (2022). Active learning (AL) addresses this challenge by iteratively selecting the most informative unlabeled samples for annotation, thereby boosting model efficiency Xie et al. (2021); Wang et al. (2023). Deep active learning strategies are typically divided into two categories: uncertainty-based and representativeness/diversity-based approaches Sener & Savarese (2017); Zhu & Bento (2017); Ash et al. (2019). Uncertainty-based methods focus on querying samples the model is most uncertain about, while diversity-based strategies aim to select a diverse subset of samples that effectively represent the entire dataset Citovsky et al. (2021). Both strategies have limitations, especially in batch selection and varying dataset complexities, leading to the inconsistency of the strategy efficiency Ren et al. (2021). Hybrid strategies, which balance uncertainty and diversity Zhan et al. (2021a; b), offer partial solutions. However, these strategies rely on strategy choices and subjective experiment settings, which can be ineffective in real-life applications.
@@
Previous works Baram et al. (2004); Hsu & Lin (2015); Pang et al. (2018) have explored selecting among different AL strategies to achieve optimal selections without human effort. SelectAL Hacohen & Weinshall (2024) introduces a method for dynamically choosing AL strategies for different deep learning tasks and budgets by estimating the relative budget size of the problem. It involves evaluating the impact of removing small subsets of data points from the unlabeled pool to predict how different AL strategies would perform. However, this strategy depends on approximating the reduction in generalization error on small subsets, which would not fully capture the complexities of real-world tasks. Zhang et al. (2024) proposes to select the optimal batch of AL strategy from hundreds of candidates, aiming to maximize future cumulative rewards based on noisy past reward observations of each candidate algorithm. However, both the computational cost and the lack of differentiability make the optimization of these works inefficient.
@@
To address the challenges mentioned above, we propose AutoAL, an automated active learning framework that leverages differentiable query strategy search. AutoAL offers two key advantages: 1) it enables automatic and efficient updates during optimization, allowing the model to adapt seamlessly; 2) it is inherently data-driven, leveraging the underlying data distribution to guide the AL strategy selection. However, integrating multiple AL strategies into the candidate pool makes achieving differentiability difficult. To overcome this issue, AutoAL leverages a bi-level optimization framework that smoothly integrates and optimizes different AL strategies. We design two neural networks: SearchNet and FitNet. FitNet is trained on the labeled dataset to adaptively yield the informativeness of each unlabeled sample. Guided by FitNet’s task loss, SearchNet is efficiently optimized to accurately select the best AL strategy. Instead of searching over a discrete set of candidate AL strategies, we relax the search space to a continuous domain, allowing SearchNet to be optimized via gradient descent based on FitNet’s output. The gradient-based optimization, which is more data-efficient than traditional black-box searches, enables AutoAL to achieve state-of-the-art performance in strategy selection with significantly reduced computational costs. Our key contributions are as follows:
@@
  • We introduce a novel AutoAL framework for differentiable AL query strategy search. To the best of our knowledge, AutoAL is the first automatic query strategy search algorithm that can be trained in a differientiable way, achieving highly efficient update of query strategy in a data-driven manner.
@@
  • AutoAL is built upon existing AL strategies (i.e., uncertainty-based and diversity-based approaches). By incorporating most AL strategies into its search space, AutoAL leverages the advanages of these strategies through finding the most effective ones for specific AL settings, objectives, or data distributions.
@@ 2 Related Work
Active Learning has been studied for decades to improve the model performance using small sets of labeled data, significantly reducing the annotation cost Cohn et al. (1996); Settles (2009); Zhan et al. (2021b). Most AL research focuses on pool-based AL, which identifies and selects the most informative samples from a large unlabeled pool iteratively. Pool-based AL sampling strategies are generally categorized into two primary branches: diversity-based and uncertainty-based. The diversity-based methods select the batch of samples that can best represent the entire dataset distribution. CoreSet Sener & Savarese (2017) selects a batch of representative points from a core set, which is a subsample of the dataset that effectively serves as a proxy for the entire set. In contrast, uncertainty-based methods focus on selecting the samples with high uncertainty, which is due to the data generation or the modeling/learning process. Bayesian Active Learning by Disagreements (BALD) Gal et al. (2017) chooses data points that are expected to maximize the mutual information between predictions and model posterior. Yoo & Kweon (2019) incorporates a loss prediction strategy by appending a compact parametric module designed to forecast the loss of unlabeled inputs relative to the target model. This is achieved by minimizing the discrepancy between the predicted loss and the actual target loss. It requires subjective judgment and strategic selection of active learning methods to be effective on specific datasets in real-world settings.
@@
Neither diversity-based nor uncertainty-based methods are perfect in all data scenarios. Diversity-based methods tend to perform better when the dataset contains rich category content and large batch size, while uncertainty-aware methods typically perform better in opposite settings Zhan et al. (2021b); Citovsky et al. (2021). Several past works have explored adaptive selection in active learning to solve this problem. A common approach is to choose the best AL strategy from a set of candidate methods and use it to query unlabeled data, such as Hacohen & Weinshall (2024) and Zhang et al. (2024) introduced in Section 1. Active Learning By Learning (ALBL) Hsu & Lin (2015) adaptively selects from a set of existing algorithms based on their estimated contribution to learning performance. It frames each candidate as a bandit machine, converting the problem into a multi-armed bandit scenario. However, none of these methods leverage the differentiable framework for automatic AL selection. Compared to the methods mentioned above, AutoAL first relaxes the search space to be continuous, thus can optimize automatically via gradient descent. This framework enables AutoAL to seamlessly integrate multiple existing AL strategies and rapidly adapt to select the optimal strategy based on the labeled pool. The simple gradient-based optimization is much more data-efficient than traditional black-box searches. Therefore, AutoAL can significantly reduce computational costs compared to these methods.
@@
Bilevel optimization is a hierarchical mathematical framework where the feasible region of one optimization task is constrained by the solution set of another optimization task Liu et al. (2021). It has gained significant attention due to its nested problem structure, which allows the two tasks to optimize jointly. Bilevel optimization adapts to many DL applications Chen et al. (2022), such as hyperparameter optimization Chen et al. (2019); MacKay et al. (2019), meta-knowledge extraction Finn et al. (2017), neural architecture search Liu et al. (2018); Hu et al. (2020), and active learning Sener & Savarese (2017). For instance, Sener & Savarese (2017) formulates AL as a bi-level coreset selection problem and designs a 2-approximation method to select a subset of data that represents the entire dataset. In this work, we novelly formulate the Al query strategy search as a bi-level optimization problem, i.e., iteratively updating FitNet and using its task loss to guide the optimization of SearchNet. We further reformulate it as a differentiable learning paradigm to enable a more efficient selection of examples across various objectives and data distributions.
@@ 3.2 Automated Active Learning
In AutoAL, each AL iteration consists of $\mathcal{C}$ cycles. In each cycle $c$, the labeled data $L$ is randomly split into two equal-sized subsets: training and validation sets. FitNet $\Omega_{F}$ computes the cross-entropy loss $\mathcal{L}$  Zhang & Sabuncu (2018) for data. Meanwhile, SearchNet $\Omega_{S}$ generates scores for the samples, which are used to rank them. This raises a key question: If only the training set is available, without access to the unlabeled pool, can $\Omega_{S}$ still effectively select the optimal samples?
Our key design is intuitive: SearchNet $\Omega_{S}$ treats the training batch as the labeled pool and the validation batch as the unlabeled pool. This allows $\Omega_{S}$ to simulate the process of AL selection, without direct access to the actual unlabeled data pool. The optimization goal of $\Omega_{S}$ is to select data with higher losses, as they are expected to provide more informative updates to the neural network. For $\Omega_{F}$, the training objective is to minimize the loss of the selected samples, particularly those prioritized by $\Omega_{S}$. AutoAL is formulated as a bi-level optimization problem, as the optimization of $\Omega_{F}$ is dependent only on itself, but the optimization of $\Omega_{S}$ depends on the optimal $\Omega_{F}$.
@@ 3.3 Differentiable Query Strategy Optimization
During training, for each labeled data $x_{j}$,we query a score $\mathcal{S}{\kappa}(x_{j})$ from each candidate AL sampling strategy $\mathcal{A}i$, where $\mathcal{S}{\kappa}(x_{j})\in{0,1}$ indicates whether the sample is selected. To model the overall score $\mathcal{S}(x_{j})$, which is a combination of all the AL search scores $\mathcal{S}{\kappa}(x_{j})$, we adopt a Gaussian Mixture approach Reynolds et al. (2009):
@@ 3.4 Learning and Algorithm of AutoAL
As described in section 3.2, in cycle $c$-2, $\Omega_{F}$ is optimized to get the data distribution. The final loss is calculated according to Eq. 7. In cycle $c$-1, AutoAL optimizes $\Omega_{S}$ to get the data sample with highest loss. The loss function from $\Omega_{F}$ is detached to avoid updating itself, and the final loss is calculated according to Eq. 8. Here we use negative loss to achieve gradient ascent. For $n\in\left\{0,1,\ldots,\left\lfloor\frac{M}{2B}\right\rfloor\right\}$, AutoAL additionally integrates a loss prediction module according to Yoo & Kweon (2019) to help update $\Omega_{S}$, as
@@
0:  $K$ candidate algorithms $\mathcal{A}=\{\mathcal{A}_{\kappa}\}_{\kappa\in[K]}$, labeled pool $L=\{(x_{j},y_{j})\}_{j=1}^{M}$ and unlabled pool $U=\{x_{i}\}_{i=1}^{N}$, total number of AL rounds $\mathcal{R}$, batch size $b$, task model $\mathcal{T}$.
0:  task model $\mathcal{T}$
1:  Initialize $\mathcal{T}$
2:  for $c$$\mathcal{R}$ do
3:     Optimize $\Omega^{*}_{F}$, $\Omega^{*}_{S}$ according to Eq. 1;
4:     Calculate $\bar{\mathcal{S}}(x_{i})$ by Eq. 6 for all $i\in N$;
5:     Select the top $b$ samples with the highest score $\bar{\mathcal{S}}(x_{j})$;
6:     Query label $y_{i}$ for all $i\in b$; Update $L$ and $U$;
7:     Train task model $\mathcal{T}$ by using $L$;
8:  end for
9:  return $\mathcal{T}$
@@ 4.1 Experiment Settings
Table 1: Overview of dataset information.
Dataset # of training set # of test set # of classes
Nature image Cifar-10 $50,000$ $10,000$ $10$
Cifar-100 $50,000$ $10,000$ $100$
SVHN $73,257$ $26,032$ $10$
Medical image OrganCMNIST $12,975$ $8,216$ $11$
PathMNIST $89,996$ $7,180$ $9$
TissueMNIST $165,466$ $47,280$ $8$
@@
We conduct AL experiments on six datasets: Cifar-10 and Cifar-100 Krizhevsky et al. (2009), SVHN Netzer et al. (2011) in the nature image domain, and OrganCMNIST, PathMNIST, and TissueMNIST from MedMNIST Yang et al. (2023) in the medical image domain. We compare our method against multiple AL sampling strategies, including Maximum Entropy Sampling Shannon (1948), Margin Sampling Netzer et al. (2011), Least Confidence Wang & Shang (2014), KMeans Sampling Ahmed et al. (2020), Bayesian Active Learning by Disagreements (BALD) Gal et al. (2017), Variation Ratios (VarRatio) Freeman (1965) and Mean Standard Deviation (MeanSTD) Kampffmeyer et al. (2016). We also consider the state-of-the-art deep AL methods as baselines, including Batch Active learning by Diverse Gradient Embeddings (BADGE) Ash et al. (2019) and Loss Prediction Active Learning (LPL) Yoo & Kweon (2019). We also include ALBL Hsu & Lin (2015) to compare AutoAL with existing selective AL strategies.
@@
For $\Omega_{F}$ and $\Omega_{S}$, we build the backbone using ResNet-18 He et al. (2016). We also employ ResNet-18 as the classification model on all baselines and AutoAL for fair comparison. Since AutoAL is built upon existing AL strategies and focuses on selecting the optimal strategy, we integrate seven AL methods into AutoAL: Maximum Entropy, Margin Sampling, Least Confidence, KMeans, BALD, VarRatio, and MeanSTD. For the optimization of $\Omega_{S}$ and loss prediction module, we use SGD optimizer Ruder (2016). For the optimization of $\Omega_{F}$, we use Adam optimizer Kingma (2014). Both use $0.005$ as the learning rate. While training, FitNet will first update for $200$ epochs using the validation queue, then $\Omega_{F}$, $\Omega_{S}$ and the loss prediction module will update iteratively with a total of $400$ epochs. All experiments are repeated three times with different randomly selected initial labeled pools, reporting mean and standard deviation.
Figure 2: Overall performance on six benchmark datasets: natural image datasets (top) and medical image datasets (bottom).

### v2

@@ Abstract
As deep learning continues to evolve, the need for data efficiency becomes increasingly important. Considering labeling large datasets is both time-consuming and expensive, active learning (AL) provides a promising solution to this challenge by iteratively selecting the most informative subsets of examples to train deep neural networks, thereby reducing the labeling cost. However, the effectiveness of different AL algorithms can vary significantly across data scenarios, and determining which AL algorithm best fits a given task remains a challenging problem. This work presents the first differentiable AL strategy search method, named AutoAL, which is designed on top of existing AL sampling strategies. AutoAL consists of two neural nets, named SearchNet and FitNet, which are optimized concurrently under a differentiable bi-level optimization framework. For any given task, SearchNet and FitNet are iteratively co-optimized using the labeled data, learning how well a set of candidate AL algorithms perform on that task. With the optimal AL strategies identified, SearchNet selects a small subset from the unlabeled pool for querying their annotations, enabling efficient training of the task model. Experimental results demonstrate that AutoAL consistently achieves superior accuracy compared to all candidate AL algorithms and other selective AL approaches, showcasing its potential for adapting and integrating multiple existing AL methods across diverse tasks and domains. Code is available at: https://github.com/haizailache999/AutoAL.
Machine Learning, ICML
@@ 1 Introduction
With the development of deep learning (DL) techniques, large volumes of high-quality labeled data have become increasingly crucial as the model training processes are usually data hungry  (Ren et al., 2021; Zhan et al., 2022). Active learning (AL) addresses this challenge by iteratively selecting the most informative unlabeled samples for annotation, thereby boosting model efficiency (Xie et al., 2021; Wang et al., 2023). Deep active learning strategies are typically divided into two categories: uncertainty-based and representativeness/diversity-based approaches (Sener & Savarese, 2017; Zhu & Bento, 2017; Ash et al., 2019). Uncertainty-based methods focus on querying samples the model is most uncertain about, while diversity-based strategies aim to select a diverse subset of samples that effectively represent the entire dataset (Citovsky et al., 2021). Both strategies have limitations, especially in batch selection and varying dataset complexities, leading to the inconsistency of the strategy efficiency (Ren et al., 2021). Hybrid strategies, which balance uncertainty and diversity (Zhan et al., 2021a, b), offer partial solutions. However, these strategies rely on strategy choices and subjective experiment settings, which can be ineffective in real-life applications.
@@
Previous works (Baram et al., 2004; Hsu & Lin, 2015; Pang et al., 2018) have explored selecting among different AL strategies to achieve optimal selections without human effort. SelectAL (Hacohen & Weinshall, 2024) introduces a method for dynamically choosing AL strategies for different deep learning tasks and budgets by estimating the relative budget size of the problem. It involves evaluating the impact of removing small subsets of data points from the unlabeled pool to predict how different AL strategies would perform. However, this strategy depends on approximating the reduction in generalization error on small subsets, which would not fully capture the complexities of real-world tasks. Zhang et al. (2024) proposes to select the optimal batch of AL strategy from hundreds of candidates, aiming to maximize future cumulative rewards based on noisy past reward observations of each candidate algorithm. However, both the computational cost and the lack of differentiability make the optimization of these works inefficient.
@@
To address these challenges, we propose AutoAL, an automated active learning framework that leverages differentiable query strategy search. AutoAL offers two key advantages: 1) it enables automatic and efficient updates during optimization, allowing the model to adapt seamlessly; 2) it is inherently data-driven, leveraging the underlying data distribution to guide the AL strategy selection. However, integrating multiple AL strategies into the candidate pool makes achieving differentiability difficult. To overcome this issue, AutoAL leverages a bi-level optimization framework that smoothly integrates and optimizes different AL strategies. We design two neural networks: SearchNet and FitNet. FitNet is trained on the labeled dataset to adaptively yield the informativeness of each unlabeled sample. Guided by FitNet’s task loss, SearchNet is efficiently optimized to accurately select the best AL strategy. Instead of searching over a discrete set of candidate AL strategies, we relax the search space to a continuous domain, allowing SearchNet to be optimized via gradient descent based on FitNet’s output. The gradient-based optimization, which is more data-efficient than traditional black-box searches, enables AutoAL to achieve state-of-the-art performance in strategy selection with significantly reduced computational costs. Our key contributions are as follows:
@@
  • We introduce a novel AutoAL framework for differentiable AL query strategy search. To the best of our knowledge, AutoAL is the first automatic query strategy search algorithm that can be trained in a differentiable way, achieving highly efficient updates of query strategy in a data-driven manner.
@@
  • AutoAL is built upon existing AL strategies (i.e., uncertainty-based and diversity-based). By incorporating most AL strategies into its search space, AutoAL leverages the advantages of these strategies by finding the most effective ones for specific AL settings, objectives, or data distributions.
@@ 2 Related Work
Active Learning has been studied for decades to improve the model performance using small sets of labeled data, significantly reducing the annotation cost (Cohn et al., 1996; Settles, 2009; Zhan et al., 2021b). Most AL research focuses on pool-based AL, which identifies and selects the most informative samples from a large unlabeled pool iteratively. Pool-based AL sampling strategies are generally categorized into two primary branches: diversity-based and uncertainty-based. The diversity-based methods select the batch of samples that can best represent the entire dataset distribution. CoreSet (Sener & Savarese, 2017) selects a batch of representative points from a core set, which is a subsample of the dataset that effectively serves as a proxy for the entire set. In contrast, uncertainty-based methods focus on selecting the samples with high uncertainty, which is due to the data generation or the modeling/learning process. Bayesian Active Learning by Disagreements (BALD) (Gal et al., 2017) chooses data points that are expected to maximize the mutual information between predictions and model posterior. Meta-Query Net (Park et al., 2022) trains an MLP that receives one open-set score and one AL score as input and outputs a balanced meta-score for sample selection in the open-set dilemma. Yoo & Kweon (2019) incorporates a loss prediction strategy by appending a compact parametric module designed to forecast the loss of unlabeled inputs relative to the target model. This is achieved by minimizing the discrepancy between the predicted loss and the actual target loss. It requires subjective judgment and strategic selection of active learning methods to be effective on specific datasets in real-world settings.
@@
Neither diversity-based nor uncertainty-based methods are perfect in all data scenarios. Diversity-based methods tend to perform better when the dataset contains rich category content and large batch size, while uncertainty-aware methods typically perform better in opposite settings (Zhan et al., 2021b; Citovsky et al., 2021). Several past works have explored adaptive selection in active learning to solve this problem. A common approach is to choose the best AL strategy from a set of candidate methods and use it to query unlabeled data, such as (Hacohen & Weinshall, 2024) and (Zhang et al., 2024) introduced in Section 1. Active Learning By Learning (ALBL) (Hsu & Lin, 2015) adaptively selects from a set of existing algorithms based on their estimated contribution to learning performance. It frames each candidate as a bandit machine, converting the problem into a multi-armed bandit scenario. However, none of these methods leverage the differentiable framework for automatic AL selection. Compared to the methods mentioned above, AutoAL first relaxes the search space to be continuous, thus can optimize automatically via gradient descent. This framework enables AutoAL to seamlessly integrate multiple existing AL strategies and rapidly adapt to select the optimal strategy based on the labeled pool. The simple gradient-based optimization is much more data-efficient than traditional black-box searches. Therefore, AutoAL can significantly reduce computational costs compared to these methods.
@@
Bilevel optimization is a hierarchical mathematical framework where the feasible region of one optimization task is constrained by the solution set of another optimization task (Liu et al., 2021). It has gained significant attention due to its nested problem structure, which allows the two tasks to optimize jointly. Bilevel optimization adapts to many DL applications (Chen et al., 2022), such as hyperparameter optimization (Chen et al., 2019; MacKay et al., 2019), meta-knowledge extraction (Finn et al., 2017), neural architecture search (Liu et al., 2018; Hu et al., 2020), and active learning (Sener & Savarese, 2017). For instance, Sener & Savarese (2017) formulates AL as a bi-level coreset selection problem and designs a 2-approximation method to select a subset of data that represents the entire dataset. In this work, we novelly formulate the Al query strategy search as a bi-level optimization problem, i.e., iteratively updating FitNet and using its task loss to guide the optimization of SearchNet. We further reformulate it as a differentiable learning paradigm to enable a more efficient selection of examples across various objectives and data distributions.
@@ 3.2 Automated Active Learning
In AutoAL, each AL iteration consists of $\mathcal{C}$ cycles. In each cycle $c$, the labeled data $L$ is randomly split into two equal-sized subsets: training and validation sets. FitNet $\Omega_{F}$ computes the cross-entropy loss $\mathcal{L}$  (Zhang & Sabuncu, 2018) for data. Meanwhile, SearchNet $\Omega_{S}$ generates scores for the samples, which are used to rank them. This raises a key question: If only the training set is available, without access to the unlabeled pool, can $\Omega_{S}$ still effectively select the optimal samples?
Our key design is intuitive: SearchNet $\Omega_{S}$ treats the training batch as the labeled pool and the validation batch as the unlabeled pool. This allows $\Omega_{S}$ to simulate the process of AL selection, without direct access to the actual unlabeled data pool. The output of $\Omega_{S}$ is shown in Figure 1, a sample-wise score that aggregates the candidate AL strategy selection. The output of $\Omega_{S}$ is the informative loss of each sample. The optimization goal of $\Omega_{S}$ is to select data with higher losses, as they are expected to provide more informative updates to the neural network. For $\Omega_{F}$, the training objective is to minimize the loss of the selected samples, particularly those prioritized by $\Omega_{S}$. AutoAL is formulated as a bi-level optimization problem, as the optimization of $\Omega_{F}$ is dependent only on itself, but the optimization of $\Omega_{S}$ depends on the optimal $\Omega_{F}$.
@@ 3.3 Differentiable Query Strategy Optimization
During training, for each labeled data $x_{j}$,we query a score $\mathcal{S}{\kappa}(x_{j})$ from each candidate AL sampling strategy $\mathcal{A}i$, where $\mathcal{S}{\kappa}(x_{j})\in{0,1}$ indicates whether the sample is selected. To model the overall score $\mathcal{S}(x_{j})$, which is a combination of all the AL search scores $\mathcal{S}{\kappa}(x_{j})$, we adopt a Gaussian Mixture approach (Reynolds et al., 2009), which has been proved efficient in AL (Zhao et al., 2020):
@@ 3.4 Learning and Algorithm of AutoAL
As described in section 3.2, in cycle $c$-2, $\Omega_{F}$ is optimized to get the data distribution. The final loss is calculated according to Eq. 7. In cycle $c$-1, AutoAL optimizes $\Omega_{S}$ to get the data sample with highest loss. The loss function from $\Omega_{F}$ is detached to avoid updating itself, and the final loss is calculated according to Eq. 8. Here we use negative loss to achieve gradient ascent. For $n\in\left\{0,1,\ldots,\left\lfloor\frac{M}{2B}\right\rfloor\right\}$, AutoAL additionally integrates a loss prediction module according to (Yoo & Kweon, 2019) to help update $\Omega_{S}$, as
@@
0:  $K$ candidate algorithms $\mathcal{A}=\{\mathcal{A}_{\kappa}\}_{\kappa\in[K]}$, labeled pool $L=\{(x_{j},y_{j})\}_{j=1}^{M}$ and unlabled pool $U=\{x_{i}\}_{i=1}^{N}$, total number of AL rounds $\mathcal{R}$, batch size $b$, task model $\mathcal{T}$. 0:  task model $\mathcal{T}$ 1:  Initialize $\mathcal{T}$ 2:  for $c$$\mathcal{R}$ do 3:    Optimize $\Omega^{*}_{F}$, $\Omega^{*}_{S}$ according to Eq. 1; 4:    Calculate $\bar{\mathcal{S}}(x_{i})$ by Eq. 6 for all $i\in N$; 5:    Select the top $b$ samples with the highest score $\bar{\mathcal{S}}(x_{j})$; 6:    Query label $y_{i}$ for all $i\in b$; Update $L$ and $U$; 7:    Train task model $\mathcal{T}$ by using $L$; 8:  end for 9:  return $\mathcal{T}$
@@ 4.1 Experiment Settings
We conduct AL experiments on seven datasets: Cifar-10 and Cifar-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011), TinyImageNet (Le & Yang, 2015) in the nature image domain, and OrganCMNIST, PathMNIST, and TissueMNIST from MedMNIST database (Yang et al., 2023) in the medical image domain. The details of datasets are presented in Appendix Table 1. We compare our method against multiple AL sampling strategies, including Maximum Entropy Sampling (Shannon, 1948), Margin Sampling (Netzer et al., 2011), Least Confidence (Wang & Shang, 2014), KMeans Sampling (Ahmed et al., 2020), Bayesian Active Learning by Disagreements (BALD) (Gal et al., 2017), Variation Ratios (VarRatio) (Freeman, 1965) and Mean Standard Deviation (MeanSTD) (Kampffmeyer et al., 2016). We also consider the state-of-the-art deep AL methods as baselines, including Batch Active learning by Diverse Gradient Embeddings (BADGE) (Ash et al., 2019), Loss Prediction Active Learning (LPL) (Yoo & Kweon, 2019), Variational Adversarial Active Learning (VAAL) (Sinha et al., 2019), Core-set Selection (Coreset) (Sener & Savarese, 2017), Ensemble Variance Ratio Learning (ENSvarR) (Beluch et al., 2018), Deep Deterministic Uncertainty (DDU) (Mukhoti et al., 2023). We also include ALBL (Hsu & Lin, 2015) to compare AutoAL with existing selective AL strategies.
@@
For $\Omega_{F}$ and $\Omega_{S}$, we build the backbone using ResNet-18 (He et al., 2016). We also employ ResNet-18 as the classification model on all baselines and AutoAL for fair comparison. Since AutoAL is built upon existing AL strategies and focuses on selecting the optimal strategy, we integrate seven AL methods into AutoAL: Maximum Entropy, Margin Sampling, Least Confidence, KMeans, BALD, VarRatio, and MeanSTD. For the optimization of $\Omega_{S}$ and loss prediction module, we use SGD optimizer (Ruder, 2016). For the optimization of $\Omega_{F}$, we use Adam optimizer (Kingma, 2014). Both use $0.005$ as the learning rate. While training, FitNet will first update for $200$ epochs using the validation queue, then $\Omega_{F}$, $\Omega_{S}$ and the loss prediction module will update iteratively with a total of $400$ epochs. All experiments are repeated three times with different randomly selected initial labeled pools, reporting mean and standard deviation.
Figure 2: Overall performance on seven benchmark datasets: natural image datasets (top) and medical image datasets (bottom).